OpenStudy (gabylovesyou):

factor p^3(p^2 - 144)

OpenStudy (anonymous):

\[a^2-b^2 \implies (a+b)(a+b)\]

OpenStudy (jdoe0001):

ding ding ding \(\bf \textit{difference of squares} \\ \quad \\ (a-b)(a+b) = a^2-b^2\qquad \qquad a^2-b^2 = (a-b)(a+b)\)

OpenStudy (anonymous):

p^5-144p^3

OpenStudy (anonymous):

batman is it correct?

OpenStudy (anonymous):

Noticed I made a mistake, should be (a+b)(a-b) [difference of squares]

OpenStudy (jdoe0001):

\(\bf p^3(p^2 - 144) \qquad {\color{brown}{ 12^2=144}}\qquad p^3(p^2 - {\color{brown}{ 12^2}})\) ding ding ding, dinner is ready is going to be a delicous "oxygen" hamburger

OpenStudy (here_to_help15):

Need help still or ....

OpenStudy (gabylovesyou):

@Here_to_Help15 not on this question.. thanks!

OpenStudy (here_to_help15):

Oh ok 0.0 oh crap just noticed i am on closed questions lol